OpenStudy (anonymous):

please help me with sigma notations:

OpenStudy (anonymous):

\[\sum_{i=3}^{8}3+6i\]

OpenStudy (anonymous):

@primeralph

OpenStudy (anonymous):

so we start at 3 and go to 8

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

but what does the equation look like?

OpenStudy (anonymous):

i'm new at this, can you help me?

OpenStudy (anonymous):

Can't we can take out 3, 6 times out of the sigma?

OpenStudy (zarkon):

(3+6(3))+(3+6(4))+...+(3+6(8))

OpenStudy (anonymous):

Firstly put i = 3, and then i = 4, and then i = 5, and then add what are you getting by putting the values.

OpenStudy (anonymous):

For i = 3, 3 + 6(3) For i = 4, 3 + 6(4) For i = 5, 3 + 6(5) ......... Like this upto 8. And add all..

OpenStudy (anonymous):

[3+6(3)]+[3+6(4)]+[3+6(5)]+[3+6(6)]+[3+6(7)]+[3+6(8)] is that what it looks like?

OpenStudy (anonymous):

You will get a number simply..

OpenStudy (primeralph):

@trumpetmaster7777 yeah

OpenStudy (anonymous):

[3+6(3)]+[3+6(4)]+[3+6(5)]+[3+6(6)]+[3+6(7)]+[3+6(8)] 3+18 3+24 3+30 3+36 3+42 3+48 21 27 33 39 45 51 21+27+33+39+45+51 =?

OpenStudy (anonymous):

Yep.. Go ahead.

OpenStudy (anonymous):

216

OpenStudy (anonymous):

is that correct?

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

Simple, no??

OpenStudy (anonymous):

You are welcome dear.

OpenStudy (anonymous):

yes, pretty easy for now.

OpenStudy (anonymous):

Keep it up then..

OpenStudy (primeralph):

medal for me?

OpenStudy (anonymous):

sry, gave it to water